A ball is thrown upward with an initial velocity of 10 m/s.
How high does the ball rise?



What is the total time flight?
1
Expert's answer
2021-09-08T16:12:58-0400

Given:

"v_0=10\:\rm m/s"


(a) the maximum height

"h_{\max}=\frac{v_0^2}{2g}=\frac{(10\:\rm m/s)^2}{2*9.8\:\rm m/s^2}=5.1\:\rm m"

(b) the total time of movement

"t=2*\frac{v_0}{g}=2*\frac{10\:\rm m/s}{9.8\:\rm m/s^2}=2.0\:\rm s"